OpenStudy (anonymous):

simplify (sqrt(x^2-x)-x)* (sqrt(x^2-x)+x/(sqrt(x^2-x)+x)

OpenStudy (lgbasallote):

\[\large (\sqrt{x^2 - x} - x )\times \frac{\sqrt{x^2 - x} + x}{\sqrt{x^2 - x} +x}\] isnt this just times 1?

OpenStudy (anonymous):

exactly, but I meant, but what does it equal at the end?!

OpenStudy (anonymous):

it's not times one. It's time the congent.

OpenStudy (lgbasallote):

\[\implies (\sqrt{x^2 - x} - x) \times 1\]

OpenStudy (lgbasallote):

did i write your question right?

OpenStudy (anonymous):

\[\large \frac{\large (\sqrt{x^2 - x} - x )\times \sqrt{x^2 - x} + x}{\sqrt{x^2 - x} +x}\]

OpenStudy (anonymous):

Although the answer will be same..

OpenStudy (anonymous):

yes, exactly, but now what? How does one solve such a problem?

OpenStudy (lgbasallote):

\[b \times 1 = b\] anything multipliedto 1 remains the same

OpenStudy (anonymous):

the answer will not be the same. It will remove the square roots

OpenStudy (anonymous):

\[\large \frac{\large (\sqrt{x^2 - x} - x )\times \cancel{\sqrt{x^2 - x} + x}}{\cancel{\sqrt{x^2 - x} +x}}\]

OpenStudy (lgbasallote):

OHHHH conjugate...

OpenStudy (anonymous):

yes! that's the word!

OpenStudy (lgbasallote):

you dont really do conjugate in this situation...but if you insist...

OpenStudy (anonymous):

if I did it wrong then tell me.

OpenStudy (anonymous):

What is your actual question ??

OpenStudy (anonymous):

the first part of the equation times the conjugate.

OpenStudy (anonymous):

\[\large (\sqrt{x^2 - x} - x )\times \frac{\sqrt{x^2 - x} + x}{\sqrt{x^2 - x} +x} \implies \frac{x^2 - x - x^2}{\sqrt{x^2 - x} +x}\]

OpenStudy (lgbasallote):

\[\frac{(\sqrt{x^2 - x} - x)(\sqrt{x^2 - x} + x)}{\sqrt{x^2 - x} + x}\] \[\implies \frac{(\sqrt{x^2 - x})^2 - (x)^2}{\sqrt{x^2 - x} + x}\] \[\implies \frac{ x^2 - x -x^2}{\sqrt{x^2 - x} + x}\] \[\implies \frac{-x}{\sqrt{x^2 - x} + x}\]

OpenStudy (anonymous):

\[\large (\sqrt{x^2 - x} - x )\times \frac{\sqrt{x^2 - x} + x}{\sqrt{x^2 - x} +x} \implies \frac{x^2 - x - x^2}{\sqrt{x^2 - x} +x} \implies \frac{-x}{\sqrt{x^2 - x} +x}\]

OpenStudy (lgbasallote):

okay i encountered some problems coz i couldnt see the live preview of the latex =)))

OpenStudy (anonymous):

Just reload the page..

OpenStudy (lgbasallote):

do you get what i did @Compgroupmail

OpenStudy (anonymous):

yes, just absorbing it now.

OpenStudy (anonymous):

alright thanks. I got it now. :)

OpenStudy (anonymous):

water, what tool do you use in displaying the equations?

OpenStudy (lgbasallote):

wonderful